# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v3

@@ Abstract
Index Terms:
Few-shot Learning, K-Nearest Neighbors, Maximum Mean Discrepancy, Cross-domain Wi-Fi Sensing, Channel State Information
@@ I Introduction
These days, the importance of wireless sensing has grown significantly, driven by advancements in technology and the increasing demand for efficient monitoring solutions. Wireless sensing has found applications in various fields such as fall detection [1], person localization [2], and people identification [3]. Among the various manners, Wi-Fi sensing stands out as a popular method with numerous advantages, as outlined in Table I. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing is privacy-preserving compared with camera. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions since its electromagnetic wave operates on a frequency band that can penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally [4]. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@
Millimeter Wave Fairly Comprehensive Quite Good Moderate Fairly Secure Expensive
@@
Wi-Fi Fairly Comprehensive Quite Good Good (non-line-of-sight, all-day) Secure Negligible
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [5], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. By using this CSI data collection tool, a series of research employing commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [6] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purposes without sensing optimization, it is thus challenging to apply physical models to infer and sense the ambient environment from the CSI measurements.
As a result, most Wi-Fi sensing methods rely on data-driven approach via machine learning [7, 8], especially for fine-grained tasks like gesture recognition [9]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
@@
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment Learning (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [11, 12, 13, 14, 15, 16, 17]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain. Another widely recognized issue in DA is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [18] and Section IV-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
@@
To address the limitations of traditional DAL methods in cross-domain Wi-Fi sensing, we propose a novel Few-Shot Learning (FSL) framework, called K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD). Our approach introduces a local distribution alignment strategy that aligns the source and target domains within each category, rather than globally aligning their overall data distributions. This refinement enhances the effectiveness of traditional DAL techniques by addressing inter-category misalignments. Additionally, our framework ensures that the support set is excluded from the network’s training process, enabling the implementation of an early stopping strategy based on support set performance. This ensures stable and consistent target domain performance while avoiding the challenges faced by traditional DAL methods in determining an optimal stopping point.
The key contributions of this work are summarized as follows:
@@
The structure of this paper is as follows. In Section II, we first analyze the cross-domain Wi-Fi sensing tasks and illustrate the limitations of traditional DA methods. In Section III, we revisit the cross-domain task and propose our local alignment method in detail. In Section IV, we use a series of experiments to demonstrate the efficiency and stability of our method. Finally, in Section V, we conclude the paper and provide some points for future research.
II Problem Statement and Existing Methods
@@ II-A Wi-Fi Sensing Principle
$H=||H||e^{j\angle H}$ (2)
where $||H||$ and $\angle H$ denote the amplitude and phase of the CSI measurement, respectively.
@@ II-B Domain Shift Issue in Wi-Fi Sensing
(a) Average
(b) Standard Deviation
Figure 2: Average and Standard Deviation of CSI Amplitude (WiGesture Dataset) [3]: Each point represents a 1 second sample. The x-axis represents different people and action IDs, while different colors represent different actions.
@@
In cross-domain Wi-Fi sensing, the most significant challenge is the substantial variation in data distribution across different domains, which causes models trained in the source domain to often fail in the target domain. For better illustration, we use WiGesture dataset [3], which contains the CSI collected from 8 individuals performing 6 different gesture actions, to illustrate the domain shift problem in detail. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains. Fig. 2 depicts the distribution of mean and standard deviation of CSI amplitude across various individuals and actions. It is evident that even for the same action, individuals exhibit distinct data distributions. Likewise, Fig. 3a and 3b display the Uniform Manifold Approximation and Projection (UMAP) based data reduction [19] results (we will introduce this data dimension reduction method in Section III-D1). It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a ResNet18 [20] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained ResNet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
Figure 3: Data Dimension Reduction Results of the WiGesture Dataset [3]: Different colors represent different categories. The data dimensions have been reduced to three for visualization, corresponding to the three axes in the figures. They remain consistent across subsequent figures.
II-C Existing Domain Adaptation Methods
@@
DAL methods aim to map the source domain and target domain to a common feature distribution, allowing the data in both domains to share the same distribution in the new space. However, most DAL methods, such as MMD [25], Kernel Mean Matching (KMM) [34], and Subspace Geodesic Flow (SGF) [35], rely on the assumption that the conditional probabilities in the source and target domains are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$, where $P_{s}$ and $P_{t}$ represent the probability distributions in the source and target domains, respectively, and $x$ and $y$ represent the data sample and ground truth category label. When training the network in the source domain, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)\approx P(y|x)$ under the source domain $P_{s}(x)$. However, since $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain $P_{t}(x)$. Consequently, these traditional methods can be viewed as attempting to find a latent space where the input data from the source and target domains are mapped to the same distribution, i.e., $P_{t}(\text{g}(x))\approx P_{s}(\text{g}(x))$, where $\text{g}(\cdot)$ is a mapping function. This allows the model to learn the parameters $\theta$ more effectively and accurately in the target domain. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not satisfied. This has been demonstrated in our experiments (Section IV-C) as well as in our previous research [18], where traditional DAL methods have even failed in cross-domain Wi-Fi sensing tasks. A more detailed derivation of the formulas can be found in Section III-B.
In conclusion, the quality of the support set makes the performance of metric-based methods unstable. For learning-based methods, they also use source domain data to train, and the influence of the support set is not as significant as for metric-based methods. However, it’s hard to identify when to stop the training process, and the accuracy in the test set always has significant fluctuations during training (shown in Section IV-C). Finally, most DAL methods do not require a support set for training, and they are not affected by the quality of the support set. But they mainly rely on the assumption of $P_{t}(y|x)=P_{s}(y|x)$, which is not always established in practice. Thus, the model would easily fail.
@@
Although some previous work has highlighted the disadvantages of conventional global alignment and proposed similar pseudo-label-based DAL methods, we want to emphasize the differences between our work and theirs.
First, we argue that some prior DAL methods still struggle in challenging scenarios. For instance, the method proposed by [36] cannot effectively address the scenario illustrated in Fig. 1, as mentioned in the previous section. Our proposed method, which directly aligns the source and target domains within each category, represents a simple yet effective approach to achieve local alignment.
The success of our method primarily relies on the quality of pseudo labels for the target domain. An inevitable issue in pseudo-label-based methods is the occurrence of incorrect classifications and noise in the data. Most previous work [37, 38, 39] first trains a network in the source domain and then uses it to classify the target domain, generating pseudo labels. Typically, these methods retain only the top classification samples from the target domain for subsequent DAL training, aiming to minimize noise and incorrect primary classifications as much as possible. While these methods have shown considerable success in fields like CV and electrocardiogram (ECG) analysis, the Wi-Fi CSI serves as a different modality, often exhibiting a completely different data distribution between source and target domains. This discrepancy can render the network trained on the source domain ineffective in the target domain. For example, using the WiGesture dataset, we train a ResNet18 in the source domain and test its performance in the target domain. The confusion matrix is shown in Fig. 4a. For category 2-4, most classification results are incorrect. Notably, in category 2, there are no correct generated labels, which hinders subsequent operations and leads to significantly low model performance. In Fig. 4b, we also display the accuracy across different categories under the top $p\%$ classification confidence level. For classes other than 0 and 4, regardless of how $p\%$ is selected, the accuracy remains low, resulting in poor quality of pseudo labels. In contrast, our method utilizes model-based KNN for pseudo label generation, which is less affected by cross-domain issues compared to neural networks. Additionally, we propose a confidence-level-based approach to mitigate the influence of noise and incorrect classifications, which has been demonstrated to be effective through sensitivity analysis.
Furthermore, our method supports an early stopping mechanism, which is not feasible for most DAL methods. This feature allows us to efficiently determine when to stop the model at the optimal point, significantly enhancing model stability.
@@ III Methodology
III-A Overview
As discussed in previous section, traditional DAL methods cannot guarantee model performance. To solve this problem, we propose a local distribution alignment approach that aligns the distribution between the source domain and the target domain within each category. (The detailed mathematical analysis will be discussed in Section III-B.) However, the category labels in the testing set cannot be accessed during training, and it is also challenging to estimate the target domain distribution solely based on the limited labeled data in the support set. As a result, we propose a preliminary strategy to first obtain some labels in the testing set to estimate the distribution.
@@ III-A Overview
III-B Local Distribution Alignment
@@
$\text{MMD}[F,p,q]:=\text{sup}_{f\in F}|\text{E}_{p}[f(x)]-\text{E}_{q}[f(x)]|\ ,$ (7)

### v4

@@ I Introduction
I-A Overview
These days, the importance of wireless sensing has grown significantly, driven by advancements in technology and the increasing demand for efficient monitoring solutions. Wireless sensing has found applications in various fields such as fall detection [2], person localization [3], and people identification [4]. Among the various manners, Wi-Fi sensing—particularly in the sub-7GHz band—stands out as a popular method with numerous advantages, as outlined in Table I. [^1] Although Wi-Fi sensing can also be implemented in the millimeter-wave band, this work focuses on sub-7GHz Wi-Fi, which is more widely deployed and better supported by existing open-source CSI extraction tools. Given that millimeter-wave Wi-Fi communication is not yet mainstream [5], all subsequent references to Wi-Fi sensing in this paper refer specifically to sub-7GHz Wi-Fi. First, Wi-Fi sensing is contactless as the object to be detected does not need to carry a device like a smartphone or smartwatch. Second, Wi-Fi sensing is privacy-preserving compared with camera. Third, Wi-Fi sensing is insensitive to occlusions and illumination conditions, as the sub-7 GHz electromagnetic waves used in Wi-Fi communication can effectively penetrate walls. Moreover, Wi-Fi sensing is ubiquitous, with over $19.5$ billion Wi-Fi devices deployed globally. With an appropriate algorithm that enables its sensing function, ubiquitous sensing function can be realized in various environments, such as homes, offices, and public spaces. Last but not least, Wi-Fi sensing has lower cost compared to dedicated radars.
@@ I-A Overview
Millimeter Wave Radar Fairly Comprehensive Quite Good Moderate Fairly Secure Expensive
@@
Sub-7 GHz Wi-Fi Fairly Comprehensive Quite Good Good (non-line-of-sight, all-day) Secure Negligible
Given the released tool that works with some specific types of commercial Network Interface Cards (NIC) [6], the Channel State Information (CSI) can be obtained, which describes how the wireless signals fade during the propagation from the sender to the receiver. Using this CSI data collection tool, a series of research employs commercial Wi-Fi NICs for different kinds of tasks. Moreover, a dedicated IEEE standard for Wi-Fi sensing named IEEE 802.11bf is approved [7] to enable faster development and practicalization of Wi-Fi sensing technologies. However, since the Orthogonal Frequency Division Multiplexing (OFDM) waveforms employed by Wi-Fi are designed for communication purposes without sensing-specific optimization—such as low-sidelobe ambiguity functions, high time–frequency resolution, or consistent pilot structures—direct range–Doppler modeling from CSI measurements is less reliable than with radar-optimized waveforms, making it challenging to apply physical models to infer the ambient environment.
I-B Cross Domain Challenge
Recently, most Wi-Fi sensing methods rely on data-driven approach via machine learning, especially for fine-grained tasks like gesture recognition [8]. However, even minor environmental changes can drastically alter the CSI patterns and significantly degrade the performance of the trained model. This environmental variability leads to a domain shift - a discrepancy between the training data distribution and real-world deployment conditions. Given the strong dependence of the training data, such domain shifts can severely undermine the generalization capability. This scenario is known as a cross-domain problem in machine learning, where the source domain (training data) and target domain (testing data) have different underlying distributions or feature spaces.
@@ I-C Existing Cross-Domain Wi-Fi Sensing Methods
As the most common method for cross-domain tasks, Domain Adaptation (DA) has been widely utilized in cross-domain Wi-Fi sensing. DA focuses on leveraging knowledge from a source domain to enhance model performance on a related but different target domain. Among these methods, Domain Alignment Learning (DAL), which aims to map the source and target domains to the same distribution, is one of the most popular approaches and has been extensively applied in Wi-Fi sensing [10, 11, 12, 13, 14, 15, 16]. However, as shown in Fig. 1, the global alignment of traditional DAL methods can lead to incorrect alignments within categories, which can hinder the effectiveness of the classification boundary trained in the source domain when applied to the target domain.
Additionally, several cross-domain Wi-Fi sensing methods focus on few-shot learning (feature learning) [17]. These methods typically follow a paradigm where a feature extractor is first trained, and then similarities of the extracted features from the source and target domains are compared for classification. Most of these methods concentrate on efficiently designing and improving the feature extractor and similarity calculation components. For instance, Ding et al. [18] propose using a CNN trained in the source domain as a feature extractor, employing cosine similarity in the feature embedding for similarity calculations. They also implement a low-pass Butterworth filter for denoising. Building on this, AutoFi [19] is introduced (without the low-pass denoising), proposing a pre-training method to enhance the feature extractor’s capacity within the neural network. In Yang et al. [20], a different similarity calculation method is employed, specifically a Siamese network [21]. They utilize a CNN combined with Bi-LSTM as the feature extractor. Expanding upon this, CrossFi [22] uses a ResNet-based Siamese architecture, introducing a template generation method to enhance feature representation and an attention-based similarity calculation method to better capture relationships between samples. In CrossFi, the application of the Siamese network is extended to a zero-shot scenario. However, a widely recognized issue in few-shot learning methods is low stability: the model’s performance on the testing set can be unstable during training (as referenced in [22] and Section V-D), with accuracy varying sharply between successive epochs. This instability suggests that simply stopping the model training at a fixed epoch may result in poor performance. If we wish to implement an early stopping method, we require labeled validation data from the target domain. However, the only labeled data available from the target domain is the support set, which is often a subset of the training set and cannot be used as a validation set.
I-D Our Contributions
@@
To address the limitations of traditional DAL methods in cross-domain Wi-Fi sensing, we propose a novel Few-Shot Learning (FSL) framework, called K-Nearest Neighbors Maximum Mean Discrepancy (KNN-MMD). Our approach introduces a local distribution alignment strategy that aligns the source and target domains within each category, rather than globally aligning their overall data distributions. This refinement enhances the effectiveness of traditional DAL techniques by addressing inter-category misalignments. Additionally, our framework ensures that the support set is excluded from the network’s training process, enabling the implementation of an early stopping strategy based on support set performance. This ensures stable and consistent target domain performance while avoiding the challenges faced by traditional DAL methods in determining an optimal stopping point. The key contributions of this work are summarized as follows:
@@
The structure of this paper is as follows. In Section II, we first introduce the Wi-Fi sensing system and analyze the cross-domain Wi-Fi sensing tasks. In Section III, we review traditional DA methods across different types and illustrates their limitations. In Section IV, we revisit the cross-domain task and propose our local alignment method in detail. In Section V, we use a series of experiments to demonstrate the efficiency and stability of our method. Finally, in Section VI, we conclude the paper and provide some points for future research.
II System Model and Problem Statement
@@ II-A Wi-Fi Sensing Principle
[…]
where $H||$ and $\angle H$ denote the amplitude and phase of the CSI measurement, respectively.
@@ II-B Domain Shift Issue in Wi-Fi Sensing
Figure 2: UMAP Dimensionality Reduction of the WiGesture Dataset [4]: In the figure, different shapes represent different person IDs, while different colors represent different action IDs. The x and y axes correspond to the two main UMAP components.
@@
In cross-domain Wi-Fi sensing, the most significant challenge is the substantial variation in data distribution across different domains, which causes models trained in the source domain to often fail in the target domain. For better illustration, we use WiGesture dataset [4], which contains the CSI collected from 8 individuals performing 6 different gesture actions, to illustrate the domain shift problem in detail. For the gesture recognition or action classification task, we take different people as different domains. And for the people identification task, we take different actions as different domains. Fig. 2 depicts the Uniform Manifold Approximation and Projection (UMAP) based data reduction [23] (this method will be introduced in Section IV-D1). It is evident that even for the same action, individuals exhibit distinct data distributions; similarly, for the same individual, different actions show varied data distributions. These significant domain gaps can pose challenges for classification models. To illustrate this more clearly, Figs. 3a and 3b display the UMAP results for single and multiple individuals. It is apparent that within a single individual, different actions exhibit significant dissimilarities. However, when considering different individuals, the boundaries separating different actions become less distinct. These findings imply that it is challenging to generalize the learned features from one set of individuals to others. In Fig. 3c, a ResNet18 [24] is trained by setting the people with IDs 1-7 as the source domain and people with ID 0 as the target domain. It can be observed that the trained ResNet18 can distinguish different actions in the source domain but fail to distinguish different actions in the target domain. This is mainly because even for the same action, different people have distinct behavioral patterns, which can lead to significant gaps in the CSI patterns between individuals. As a result, to make the model perform well in the target domain, it is necessary to address the domain shift (cross-domain) problem.
@@
Figure 3: Data Dimension Reduction Results of the WiGesture Dataset [4]: Different colors represent different action categories. The data dimensions have been reduced to three for visualization, corresponding to the three axes in the figures. They remain consistent across subsequent figures.
III Preliminary
@@
DAL methods aim to map the source domain and target domain to a common feature distribution, allowing the data in both domains to share the same distribution in the new space. However, most DAL methods, such as MMD [28], Kernel Mean Matching (KMM) [37], and Subspace Geodesic Flow (SGF) [38], rely on the assumption that the conditional probabilities in the source and target domains are the same, i.e., $P_{t}(y|x)=P_{s}(y|x)=P(y|x)$, where $P_{s}$ and $P_{t}$ represent the probability distributions in the source and target domains, respectively, and $x$ and $y$ represent the data sample and ground truth category label. When training the network in the source domain, the network parameters $\theta$ need to be learned to achieve $P(y|x;\theta)\approx P(y|x)$ under the source domain $P_{s}(x)$. However, since $P_{t}(x)\neq P_{s}(x)$, the $P(y|x;\theta)$ learned from the source domain cannot generalize to the target domain $P_{t}(x)$. Consequently, these traditional methods can be viewed as attempting to find a latent space where the input data from the source and target domains are mapped to the same distribution, i.e., $P_{t}(\text{g}(x))\approx P_{s}(\text{g}(x))$, where $\text{g}(\cdot)$ is a mapping function. This allows the model to learn the parameters $\theta$ more effectively and accurately in the target domain. However, in many practical scenarios, the assumption $P_{t}(y|x)=P_{s}(y|x)$ is not satisfied. This has been demonstrated in our experiments (Section V-C) as well as in our previous research [22], where traditional DAL methods have even failed in cross-domain Wi-Fi sensing tasks. A more detailed derivation of the formulas can be found in Section IV-B.
In conclusion, the quality of the support set makes the performance of metric-based methods unstable. For learning-based methods, they also use source domain data to train, and the influence of the support set is not as significant as for metric-based methods. However, it’s hard to identify when to stop the training process, and the accuracy in the test set always has significant fluctuations during training (shown in Section V-C). Finally, most DAL methods do not require a support set for training, and they are not affected by the quality of the support set. But they mainly rely on the assumption of $P_{t}(y|x)=P_{s}(y|x)$, which is not always established in practice. Thus, the model would easily fail.
@@
Although some previous work has highlighted the disadvantages of conventional global alignment and proposed similar pseudo-label-based DAL methods, we want to emphasize the differences between our work and theirs. First, we argue that some prior DAL methods still struggle in challenging scenarios. For instance, the method proposed by [39] cannot effectively address the scenario illustrated in Fig. 1, as mentioned in the previous section. Our proposed method, which directly aligns the source and target domains within each category, represents a simple yet effective approach to achieve local alignment. The success of our method primarily relies on the quality of pseudo labels for the target domain. An inevitable issue in pseudo-label-based methods is the occurrence of incorrect classifications and noise in the data. Most previous work [40, 41, 42] first trains a network in the source domain and then uses it to classify the target domain, generating pseudo labels. Typically, these methods retain only the top classification samples from the target domain for subsequent DAL training, aiming to minimize noise and incorrect primary classifications as much as possible. While these methods have shown considerable success in fields like CV and electrocardiogram (ECG) analysis, the Wi-Fi CSI serves as a different modality, often exhibiting a completely different data distribution between source and target domains. This discrepancy can render the network trained on the source domain ineffective in the target domain. For example, using the WiGesture dataset, we train a ResNet18 in the source domain and test its performance in the target domain. The confusion matrix is shown in Fig. 4a. For category 2-4, most classification results are incorrect. Notably, in category 2, there are no correct generated labels, which hinders subsequent operations and leads to significantly low model performance. In Fig. 4b, we also display the accuracy across different categories under the top $p\%$ classification confidence level. For classes other than 0 and 4, regardless of how $p\%$ is selected, the accuracy remains low, resulting in poor quality of pseudo labels. In contrast, our method utilizes model-based KNN for pseudo label generation, which is less affected by cross-domain issues compared to neural networks. Additionally, we propose a confidence-level-based approach to mitigate the influence of noise and incorrect classifications, which has been demonstrated to be effective through sensitivity analysis. Furthermore, our method supports an early stopping mechanism, which is not feasible for most DAL methods. This feature allows us to efficiently determine when to stop the model at the optimal point, significantly enhancing model stability.
@@ IV Methodology
IV-A Overview
As discussed in previous section, traditional DAL methods cannot guarantee model performance. To solve this problem, we propose a local distribution alignment approach that aligns the distribution between the source domain and the target domain within each category. (The detailed mathematical analysis will be discussed in Section IV-B.) However, the category labels in the testing set cannot be accessed during training, and it is also challenging to estimate the target domain distribution solely based on the limited labeled data in the support set. As a result, we propose a preliminary strategy to first obtain some labels in the testing set to estimate the distribution.
@@ IV-A Overview
IV-B Local Distribution Alignment
@@
$\text{MMD}[F,p,q]=\text{sup}_{f\in F}|\text{E}_{p}[f(x)]-\text{E}_{q}[f(x)]|\ ,$ (7)